# SG-I2V: Self-Guided Trajectory Control in Image-to-Video Generation

### v1

@@ Abstract
Methods for image-to-video generation have achieved impressive, photo-realistic quality. However, adjusting specific elements in generated videos, such as object motion or camera movement, is often a tedious process of trial and error, e.g., involving re-generating videos with different random seeds. Recent techniques address this issue by fine-tuning a pre-trained model to follow conditioning signals, such as bounding boxes or point trajectories. Yet, this fine-tuning procedure can be computationally expensive, and it requires datasets with annotated object motion, which can be difficult to procure. In this work, we introduce SG-I2V, a framework for controllable image-to-video generation that is self-guided—offering zero-shot control by relying solely on the knowledge present in a pre-trained image-to-video diffusion model without the need for fine-tuning or external knowledge. Our zero-shot method outperforms unsupervised baselines while being competitive with supervised models in terms of visual quality and motion fidelity. Additional details and video results are available on our project page: https://kmcode1.github.io/Projects/SG-I2V.
@@ 1 Introduction
Recent advances in video diffusion models demonstrate significant improvements in visual and motion quality  (Ho et al., 2022b; Blattmann et al., 2023a; b; He et al., 2022). These models typically take a text prompt (Ho et al., 2022b; Blattmann et al., 2023a; Ho et al., 2022a) or image (Chen et al., 2023; 2024a; Guo et al., 2024; Xing et al., 2024) as input and generate video frames of a photorealistic, animated scene. Current methods can generate videos that are largely consistent with an input text description or image; however, fine-grained adjustment of specific video elements (e.g., object motion or camera movement) is conventionally a tedious process that requires re-running the model with different text prompts or random seeds (Wu et al., 2024a; Qiu et al., 2024).
Approaches for controllable video generation aim to eliminate this process of trial-and-error through direct manipulation of generated video elements, such as object motion (Wu et al., 2024b; Yin et al., 2023; Wang et al., 2024a), pose (Hu, 2024; Xu et al., 2024b), and camera movement (Wang et al., 2024c; Li et al., 2024; He et al., 2024a; Hu et al., 2024). One line of work fine-tunes pre-trained video generators to incorporate control signals such as bounding boxes or point trajectories (Wu et al., 2024b; Wang et al., 2024c). One of the primary challenges with these supervised methods is the expensive training cost, and thus, previous methods usually incorporate trajectory control by fine-tuning at a lower resolution than the original model Wu et al. (2024b); Yin et al. (2023). More recently, several methods for zero-shot, controllable text-to-video generation have been developed (Ma et al., 2023; Qiu et al., 2024; Jain et al., 2024). They control object trajectories by modulating the cross-attention maps between features within a bounding box and an object-related text token. Still, it is not always possible to associate a desired edit with the input text prompt (consider, e.g., motion of object parts). Moreover, these methods cannot be directly applied to animate existing images, as they are only conditioned on text.
@@
  • We conduct extensive experiments to show superior performance over zero-shot baselines while being competitive with supervised baselines in visual and motion quality.
@@ 2 Related Work
With the recent advances in diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), image animation has achieved tremendous progress (Wang et al., 2023; Chen et al., 2023; 2024a; Guo et al., 2024). Early methods inflate pre-trained text-to-image models (Rombach et al., 2022) to add motions to an image (Wu et al., 2023; Khachatryan et al., 2023; Singer et al., 2023). A notable example is AnimateDiff (Guo et al., 2024), which learns low-rank adapters (Hu et al., 2022) for different motions. Later works seek to inject a conditioning frame into a pre-trained text-to-video model (Ho et al., 2022b; a; He et al., 2022). VideoCrafter1 (Chen et al., 2023) leverages a dual cross-attention layer to condition on features of both the image and the text prompt. DynamicCrafter (Xing et al., 2024) further improves it by concatenating the input image with noisy latent. Stable Video Diffusion (SVD) (Blattmann et al., 2023a) instead works in an image-only manner, which replaces the CLIP (Radford et al., 2021) text embedding of the text prompt with the CLIP image embedding of the conditioning frame. However, none of these models support direct trajectory control of scene elements, and rather require multiple attempts to obtain a desired result. In this work, we aim to enable intuitive motion control in animating a pre-existing image. Since SVD only takes in an image without any text prompt, we utilize it as the base model following prior work (Yin et al., 2023; Wu et al., 2024b).
@@
One common way to incorporate spatial control into the image generation process is to fine-tune pre-trained models to incorporate conditioning on depth maps or bounding boxes (Zhang et al., 2023; Ye et al., 2023; Avrahami et al., 2023; Li et al., 2023; Goel et al., 2024; Wang et al., 2024b). While these methods demonstrate high fidelity, they require excessive computing resources and labor-intensive data annotations. Therefore, several tuning-free approaches have been proposed (Cao et al., 2023; Chen et al., 2024b; Feng et al., 2023; Hertz et al., 2023). Self-Guidance (Epstein et al., 2023) and Attend-and-Excite (Chefer et al., 2023) control the image layout by manipulating intermediate attention maps. They first estimate the generated objects’ positions using the attention maps produced from text–image cross-attention layers and then optimize the latent to increase the attention values at specific positions. However, these approaches require associating the control target with a specific token in the text prompt and thus cannot be directly applied to image-only editing. Closest to ours are methods that alter image layouts without text input (Pan et al., 2023; Mou et al., 2024a; b; Ling et al., 2023; Liu et al., 2024; Zhang et al., 2024b; Cui et al., 2024; Hou et al., 2024b; Zhao et al., 2024; Shi et al., 2024). They enforce similarity between semantically correlated feature maps extracted from the upsampling blocks of a denoising U-Net (Tang et al., 2023; Namekata et al., 2024; Hedlin et al., 2024; Zhang et al., 2024a; Luo et al., 2023). In contrast, we show that the cross-frame semantic correspondence of upsampling feature maps in image-to-video diffusion models (Blattmann et al., 2023a) is weak. Thus, optimization directly based on these feature maps leads to sub-optimal results.
@@
Several recent works have studied camera pose control in video diffusion models (He et al., 2024a; Xu et al., 2024a; Kuang et al., 2024; Hu et al., 2024; Xiao et al., 2024; Hou et al., 2024a; Bahmani et al., 2024; Li et al., 2024). The representative work MotionCtrl (Wang et al., 2024c) fine-tunes pre-trained video generators to follow camera trajectory input. However, video datasets with accurate camera pose annotations are limited (Zhou et al., 2018), and fine-tuning video models requires high computation costs. Another line of work focuses on object trajectory control (Wang et al., 2023; Yin et al., 2023; Wu et al., 2024b; Zhou et al., 2024; Wang et al., 2024a; Zhang et al., 2024c), among which our work is particularly related to the tuning-free variants (Qiu et al., 2024; Ma et al., 2023; Yang et al., 2024; Jain et al., 2024; Yu et al., 2024). Yet, all of these methods focus on text-based generation. In contrast, our method uses an image-to-video model and thus can turn existing, real-world images into controllable videos. Moreover, we can control camera motion by specifying trajectories of background regions.
@@ 3.1 Trajectory Control in Image-to-Video Generation
The goal of our work is to build a zero-shot framework that takes an input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ (with height $H$, width $W$) and generates a video with $N$ frames, where elements in the input image move in a user-specified fashion. Inspired by DragAnything (Wu et al., 2024b), we further assume that a user provides a set of $B$ input bounding box trajectories $\{\mathcal{B}_{b}\}_{b=1}^{B}$, each parameterized by a height $h_{b}$ and width $w_{b}$, as well as center point coordinates for each output video frame: $\mathbf{c}_{b,n}\in\mathbb{R}^{2},1\leq n\leq N$. For simplicity, we assume bounding boxes cannot extend outside the image, and we denote the $b$-th bounding box in the $n$-th frame as $\mathcal{B}_{b,n}=\{h_{b},w_{b},\mathbf{c}_{b,n}\}$.
@@ 3.2 Extracting Semantic Video Layout
Feature alignment with modified self-attention. SVD leverages separate spatial and temporal self-attention to model the entire video. Since spatial self-attention is only applied per frame, it does not produce cross-frame aligned features. While temporal attention communicates across frames, it only attends to the same pixel position on the feature map, which may be inadequate for capturing semantic information spatially. To address this issue, inspired by (Wu et al., 2023), we modify the spatial self-attention on each frame to directly attend to the first frame. Concretely, for the $n$-th frame, the original spatial self-attention works as $\bm{F}_{n}=\mathrm{Softmax}(\frac{\bm{Q}_{n}\cdot\bm{K}_{n}^{T}}{\sqrt{D}})\cdot\bm{V}_{n}$, where $\bm{F}_{n}$ is the outputs of self-attention, $\bm{Q},\bm{K},\bm{V}$ are the query, key, and value tokens, respectively, and $D$ is the dimensionality of the key and query tokens Vaswani et al. (2017). Instead, we replace the key $\bm{K}_{n}$ and value $\bm{V}_{n}$ of each frame with $\bm{K}_{1}$ and $\bm{V}_{1}$ from the first frame, leading to a new operation $\tilde{\bm{F}}_{n}=\mathrm{SoftMax}(\frac{\bm{Q}_{n}\cdot\mathrm{SG}(\bm{K}_{1})^{T}}{\sqrt{D}})\cdot\mathrm{SG}(\bm{V}_{1})$. Notably, we apply a stop gradient $\mathrm{SG}(\cdot)$ on $\bm{K}_{1}$ and $\bm{V}_{1}$ to stabilize the subsequent optimization process. Now, all the modified feature maps $\tilde{\bm{F}}_{n}$ are weighted combinations of $\bm{V}_{1}$, exhibiting a stronger cross-frame correspondence while still maintaining the object layout of each frame, as shown in the bottom row of Fig. 2.
@@ 3.3 Trajectory Control with Latent Optimization
So far, we have obtained spatio-temporally aligned feature maps $\tilde{\bm{F}}_{n}\in\mathbb{R}^{h\times w\times d}$ at each frame. For simplicity, we resize $\tilde{\bm{F}}_{n}$ extracted from different U-Net stages to the same resolution as the noisy latent $\bm{z}_{t}$. Recall that our goal is to control the output video frames so that the bounding boxes $\mathcal{B}_{b}$ identified in the first frame move along the associated trajectories. Inspired by prior drag-based control methods (Shi et al., 2024; Pan et al., 2023), we optimize the noisy latent $\bm{z}_{t}$ to enforce cross-frame similarity between features within bounding boxes. The optimization objective is as follows:
@@
where $\odot$ indicates element-wise multiplication, and we optimize the latent code $\bm{z}_{t}$ at the denoising timestep $t$. Following DragAnything (Wu et al., 2024b), we weigh the feature difference using a Gaussian heatmap $\bm{G}_{b}\in\mathbb{R}^{h\times w}$, which applies a stronger weight to the bounding box center.
@@ 3.4 High-Frequency Preserved Post-Processing
Although the presented pipeline already enables trajectory control in video generation, we notice a quality degradation in generated videos. We attribute this degradation to the deviation of $\bm{z}_{t}^{*}$ from the sampling distribution of the diffusion process after optimization. A recent work FreeInit (Wu et al., 2024a) observed that motions of generated videos are mostly encoded in the low-frequency component of noisy latent. Inspired by this, we propose to discard the high-frequency component of the optimized latent $\bm{z}_{t}^{*}$ and replace it with the high-frequency component of the original latent $\bm{z}_{t}$. Formally, we obtain the new latent $\tilde{\bm{z}}_{t}$ as follows:
$\tilde{\bm{z}}_{t}=\textsc{IFFT}_{\mathrm{2D}}\left(\textsc{FFT}_{\mathrm{2D}}(\bm{z}_{t}^{*})\odot\mathbf{H}_{\gamma}+\textsc{FFT}_{\mathrm{2D}}(\bm{z}_{t}^{*})\odot(1-\mathbf{H}_{\gamma})\right),$ (2)
@@ 4.1 Experimental Setup
In all experiments, we leverage the image-to-video variant of Stable Video Diffusion (Blattmann et al., 2023a) to generate videos with 14 frames and $576\times 1024$ resolution. The default discrete Euler scheduler (Karras et al., 2022) is applied with $T=50$ sampling steps. We extract feature maps from the last two self-attention layers from the middle stage in the denoising U-Net. We optimize Eq. 1 at the early denoising timesteps $t\in[45,44,...,30]$ for $5$ iterations per timestep. We use the AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of $0.21$. All these design choices are carefully ablated and analyzed in Sec. 4.4. During loss calculation, Gaussian heatmap $\bm{G}_{b}$ is constructed following (Wu et al., 2024b), where a heatmap for a bounding box of size $(h_{b},w_{b})$ is created by Gaussian distribution with standard deviation $\sigma=(0.2h_{b},0.2w_{b})$. For the low-pass filter $\mathbf{H}_{\gamma}$, we set the cut-off frequency $\gamma$ to $0.5$.
@@
We compare with methods that enable motion control on existing images and have publicly available code. Specifically, we compare with supervised baselines DragNUWA (Yin et al., 2023) and DragAnything (Wu et al., 2024b), which both add motion adapters to SVD. We also compare with Image Conductor (Li et al., 2024), which is based on AnimateDiff (Guo et al., 2024). Since all supervised baselines are fine-tuned to generate videos at a lower resolution, we resize the generated videos from all the methods to $320\times 576$ for a fair comparison. Since no previous methods exist for zero-shot trajectory-controlled image-to-video generation, we adapt techniques from text-to-video methods to create new baselines. Specifically, FreeTraj incorporates the noise initialization technique from (Qiu et al., 2024) by copy-pasting the initial noise on the first frame to other frames along the trajectories. We cannot apply the attention manipulation technique in their work as it is primarily designed for single-object control. DragDiffusion is inspired by the image editing method in (Shi et al., 2024), which utilizes feature maps extracted from the output of upsampling blocks to guide the generation process without feature alignment.
@@
Following prior works (Wu et al., 2024b; Zhou et al., 2024), we evaluate our method on the validation set of the VIPSeg dataset (Miao et al., 2022). We test on the same control regions and target trajectories as DragAnything, where the size of our bounding boxes is the same as the diameter of the circles in their work. For quantitative metrics, we report Frechet Inception Distance (FID) (Heusel et al., 2017) and Frechet Video Distance (FVD) (Unterthiner et al., 2018) to measure the visual quality, and ObjMC (Wu et al., 2024b) to measure the motion fidelity. ObjMC computes the average distance between generated and target trajectories, where Co-Tracker (Karaev et al., 2024) is used to estimate the trajectory of generated videos.

### v2

@@ Abstract
Methods for image-to-video generation have achieved impressive, photo-realistic quality. However, adjusting specific elements in generated videos, such as object motion or camera movement, is often a tedious process of trial and error, e.g., involving re-generating videos with different random seeds. Recent techniques address this issue by fine-tuning a pre-trained model to follow conditioning signals, such as bounding boxes or point trajectories. Yet, this fine-tuning procedure can be computationally expensive, and it requires datasets with annotated object motion, which can be difficult to procure. In this work, we introduce SG-I2V, a framework for controllable image-to-video generation that is self-guided—offering zero-shot control by relying solely on the knowledge present in a pre-trained image-to-video diffusion model without the need for fine-tuning or external knowledge. Our zero-shot method outperforms unsupervised baselines while significantly narrowing down the performance gap with supervised models in terms of visual quality and motion fidelity. Additional details and video results are available on our project page: https://kmcode1.github.io/Projects/SG-I2V.
@@ 1 Introduction
Recent advances in video diffusion models demonstrate significant improvements in visual and motion quality  (Ho et al., 2022b; Blattmann et al., 2023a; b; He et al., 2022). These models typically take a text prompt (Ho et al., 2022b; Blattmann et al., 2023a; Ho et al., 2022a) or image (Chen et al., 2023; 2024a; Guo et al., 2024; Xing et al., 2024) as input and generate video frames of a photorealistic, animated scene. Current methods can generate videos that are largely consistent with an input text description or image; however, fine-grained adjustment of specific video elements (e.g., object motion or camera movement) is conventionally a tedious process that requires re-running the model with different text prompts or random seeds (Wu et al., 2024b; Qiu et al., 2024).
Approaches for controllable video generation aim to eliminate this process of trial-and-error through direct manipulation of generated video elements, such as object motion (Wu et al., 2024c; Yin et al., 2023; Wang et al., 2024a), pose (Hu, 2024; Xu et al., 2024b), and camera movement (Wang et al., 2024c; Li et al., 2024; He et al., 2024a; Hu et al., 2024). One line of work fine-tunes pre-trained video generators to incorporate control signals such as bounding boxes or point trajectories (Wu et al., 2024c; Wang et al., 2024c). One of the primary challenges with these supervised methods is the expensive training cost, and thus, previous methods usually incorporate trajectory control by fine-tuning at a lower resolution than the original model Wu et al. (2024c); Yin et al. (2023). More recently, several methods for zero-shot, controllable text-to-video generation have been developed (Ma et al., 2023; Qiu et al., 2024; Jain et al., 2024). They control object trajectories by modulating the cross-attention maps between features within a bounding box and an object-related text token. Still, it is not always possible to associate a desired edit with the input text prompt (consider, e.g., motion of object parts). Moreover, these methods cannot be directly applied to animate existing images, as they are only conditioned on text.
@@
  • We conduct extensive experiments to show superior performance over zero-shot baselines while significantly narrowing down the performance gap with supervised baselines in visual and motion quality.
@@ 2 Related Work
With the recent advances in diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), image animation has achieved tremendous progress (Wang et al., 2023; Chen et al., 2023; 2024a; Guo et al., 2024). Early methods inflate pre-trained text-to-image models (Rombach et al., 2022) to add motions to an image (Wu et al., 2023; Khachatryan et al., 2023; Singer et al., 2023). A notable example is AnimateDiff (Guo et al., 2024), which learns low-rank adapters (Hu et al., 2022) for different motions. Later works seek to inject a conditioning frame into a pre-trained text-to-video model (Ho et al., 2022b; a; He et al., 2022). VideoCrafter1 (Chen et al., 2023) leverages a dual cross-attention layer to condition on features of both the image and the text prompt. DynamicCrafter (Xing et al., 2024) further improves it by concatenating the input image with noisy latent. Stable Video Diffusion (SVD) (Blattmann et al., 2023a) instead works in an image-only manner, which replaces the CLIP (Radford et al., 2021) text embedding of the text prompt with the CLIP image embedding of the conditioning frame. However, none of these models support direct trajectory control of scene elements, and rather require multiple attempts to obtain a desired result. In this work, we aim to enable intuitive motion control in animating a pre-existing image. Since SVD only takes in an image without any text prompt, we utilize it as the base model following prior work (Yin et al., 2023; Wu et al., 2024c).
@@
One common way to incorporate spatial control into the image generation process is to fine-tune pre-trained models to incorporate conditioning on depth maps or bounding boxes (Zhang et al., 2023; Ye et al., 2023; Avrahami et al., 2023; Li et al., 2023; Goel et al., 2024; Wang et al., 2024b). While these methods demonstrate high fidelity, they require excessive computing resources and labor-intensive data annotations. Therefore, several tuning-free approaches have been proposed (Cao et al., 2023; Chen et al., 2024b; Feng et al., 2023; Hertz et al., 2023). Self-Guidance (Epstein et al., 2023) Attend-and-Excite (Chefer et al., 2023) , and TraDiffusion (Wu et al., 2024a) control the image layout by manipulating intermediate attention maps. They first estimate the generated objects’ positions using the attention maps produced from text–image cross-attention layers and then optimize the latent to increase the attention values at specific positions. However, these approaches require associating the control target with a specific token in the text prompt and thus cannot be directly applied to image-only editing. Closest to ours are methods that alter image layouts without text input (Pan et al., 2023; Mou et al., 2024a; b; Ling et al., 2023; Liu et al., 2024; Zhang et al., 2024c; Cui et al., 2024; Hou et al., 2024b; Zhao et al., 2024; Shi et al., 2024). They enforce similarity between semantically correlated feature maps extracted from the upsampling blocks of a denoising U-Net (Tang et al., 2023; Namekata et al., 2024; Hedlin et al., 2024; Zhang et al., 2024a; Luo et al., 2023). In contrast, we show that the cross-frame semantic correspondence of upsampling feature maps in image-to-video diffusion models (Blattmann et al., 2023a) is weak. Thus, optimization directly based on these feature maps leads to sub-optimal results.
@@
Several recent works have studied camera pose control in video diffusion models (He et al., 2024a; Xu et al., 2024a; Kuang et al., 2024; Hu et al., 2024; Xiao et al., 2024b; Hou et al., 2024a; Bahmani et al., 2024; Li et al., 2024; Zhang et al., 2024b). The representative work MotionCtrl (Wang et al., 2024c) fine-tunes pre-trained video generators to follow camera trajectory input. However, video datasets with accurate camera pose annotations are limited (Zhou et al., 2018), and fine-tuning video models requires high computation costs. Another line of work focuses on object trajectory control (Wang et al., 2023; Yin et al., 2023; Wu et al., 2024c; Zhou et al., 2024; Wang et al., 2024a; Zhang et al., 2024d), among which our work is particularly related to the tuning-free variants (Qiu et al., 2024; Ma et al., 2023; Yang et al., 2024; Jain et al., 2024; Yu et al., 2024). Yet, all of these methods focus on text-based generation. In contrast, our method uses an image-to-video model and thus can turn existing, real-world images into controllable videos. Moreover, we can control camera motion by specifying trajectories of background regions.
@@ 3.1 Trajectory Control in Image-to-Video Generation
The goal of our work is to build a zero-shot framework that takes an input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ (with height $H$, width $W$) and generates a video with $N$ frames, where elements in the input image move in a user-specified fashion. Inspired by DragAnything (Wu et al., 2024c), we further assume that a user provides a set of $B$ input bounding box trajectories $\{\mathcal{B}_{b}\}_{b=1}^{B}$, each parameterized by a height $h_{b}$ and width $w_{b}$, as well as center point coordinates for each output video frame: $\mathbf{c}_{b,n}\in\mathbb{R}^{2},1\leq n\leq N$. For simplicity, we assume bounding boxes cannot extend outside the image, and we denote the $b$-th bounding box in the $n$-th frame as $\mathcal{B}_{b,n}=\{h_{b},w_{b},\mathbf{c}_{b,n}\}$.
@@ 3.2 Extracting Semantic Video Layout
Feature alignment with modified self-attention. SVD leverages separate spatial and temporal self-attention to model the entire video. Since spatial self-attention is only applied per frame, it does not produce cross-frame aligned features. While temporal attention communicates across frames, it only attends to the same pixel position on the feature map, which may be inadequate for capturing semantic information spatially. To address this issue, inspired by (Wu et al., 2023), we modify the spatial self-attention on each frame to directly attend to the first frame. Concretely, for the $n$-th frame, the original spatial self-attention works as $\bm{F}_{n}=\mathrm{Softmax}(\frac{\bm{Q}_{n}\cdot\bm{K}_{n}^{T}}{\sqrt{D}})\cdot\bm{V}_{n}$, where $\bm{F}_{n}$ is the outputs of self-attention, $\bm{Q},\bm{K},\bm{V}$ are the query, key, and value tokens, respectively, and $D$ is the dimensionality of the key and query tokens Vaswani et al. (2017). Instead, we replace the key $\bm{K}_{n}$ and value $\bm{V}_{n}$ of each frame with $\bm{K}_{1}$ and $\bm{V}_{1}$ from the first frame, leading to a new operation $\tilde{\bm{F}}_{n}=\mathrm{SoftMax}(\frac{\bm{Q}_{n}\cdot\mathrm{SG}(\bm{K}_{1})^{T}}{\sqrt{D}})\cdot\mathrm{SG}(\bm{V}_{1})$. We apply a stop gradient $\mathrm{SG}(\cdot)$ on $\bm{K}_{1}$ and $\bm{V}_{1}$ to stabilize the subsequent optimization process. Now, all the modified feature maps $\tilde{\bm{F}}_{n}$ are weighted combinations of $\bm{V}_{1}$, exhibiting a stronger cross-frame correspondence while still maintaining the object layout of each frame, as shown in the bottom row of Fig. 2. Notably, this modification occurs during the loss computation only and does not affect the denoising steps.
@@ 3.3 Trajectory Control with Latent Optimization
So far, we have obtained spatio-temporally aligned feature maps $\tilde{\bm{F}}_{n}(z_{t})\in\mathbb{R}^{h\times w\times d}$ at each frame given the noisy latent $\bm{z}_{t}$ as input (for simplicity, we resize $\tilde{\bm{F}}_{n}$ to the same resolution as the noisy latent). Recall that our goal is to control the output video frames so that the bounding boxes $\mathcal{B}_{b}$ identified in the first frame move along the associated trajectories. Inspired by prior drag-based control methods (Shi et al., 2024; Pan et al., 2023), we optimize the noisy latent $\bm{z}_{t}$ to enforce cross-frame similarity between features within bounding boxes. The optimization objective is as follows:
@@
where $\odot$ is Hadamard product, and $\tilde{\bm{F}}_{n}(\bm{z}_{t})[\mathcal{B}_{b,n}]\in\mathbb{R}^{h_{b}\times w_{b}\times d}$ is feature maps cropped by the bounding box $\mathcal{B}_{b,n}$. Following DragAnything (Wu et al., 2024c), we weigh the feature difference using a Gaussian heatmap $\bm{G}_{b}\in\mathbb{R}^{h\times w}$. This focuses on optimizing pixels closer to the bounding box center, as pixels near the edge may be background pixels that we do not want to move.
@@ 3.4 High-Frequency Preserved Post-Processing
Although the presented pipeline already enables trajectory control in video generation, we notice a quality degradation in generated videos. We attribute this degradation to the deviation of $\bm{z}_{t}^{*}$ from the sampling distribution of the diffusion process after optimization. A recent work FreeInit (Wu et al., 2024b) observed that motions of generated videos are mostly encoded in the low-frequency component of noisy latent. Inspired by this, we propose to discard the high-frequency component of the optimized latent $\bm{z}_{t}^{*}$ and replace it with the high-frequency component of the original latent $\bm{z}_{t}$. Formally, we obtain the new latent $\tilde{\bm{z}}_{t}$ as follows:
$\tilde{\bm{z}}_{t}=\textsc{IFFT}_{\mathrm{2D}}\left(\textsc{FFT}_{\mathrm{2D}}(\bm{z}_{t}^{*})\odot\mathbf{H}_{\gamma}+\textsc{FFT}_{\mathrm{2D}}(\bm{z}_{t})\odot(1-\mathbf{H}_{\gamma})\right),$ (2)
@@ 4.1 Experimental Setup
In all experiments, we leverage the image-to-video variant of Stable Video Diffusion (Blattmann et al., 2023a) to generate videos with 14 frames and $576\times 1024$ resolution. The default discrete Euler scheduler (Karras et al., 2022) is applied with $T=50$ sampling steps. We extract feature maps from the last two self-attention layers from the middle stage in the denoising U-Net. We optimize Eq. 1 at the early denoising timesteps $t\in[45,44,...,30]$ for $5$ iterations per timestep. We use the AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of $0.21$. All these design choices are carefully ablated and analyzed in Sec. 4.4. During loss calculation, Gaussian heatmap $\bm{G}_{b}$ is constructed following (Wu et al., 2024c), where a heatmap for a bounding box of size $(h_{b},w_{b})$ is created by Gaussian distribution with standard deviation $\sigma=(0.2h_{b},0.2w_{b})$. For the low-pass filter $\mathbf{H}_{\gamma}$, we set the cut-off frequency $\gamma$ to $0.5$.
@@
We compare with methods that enable motion control on existing images and have publicly available code. Specifically, we compare with supervised baselines DragNUWA (Yin et al., 2023) and DragAnything (Wu et al., 2024c), which both add motion adapters to SVD. We also compare with Image Conductor (Li et al., 2024), which is based on AnimateDiff (Guo et al., 2024). Since all supervised baselines are fine-tuned to generate videos at a lower resolution, we resize the generated videos from all the methods to $320\times 576$ for a fair comparison. Since no previous methods exist for zero-shot trajectory-controlled image-to-video generation, we adopt techniques from text-to-video methods to create new baselines. Specifically, FreeTraj incorporates the noise initialization technique from (Qiu et al., 2024) by copy-pasting the initial noise on the first frame to other frames along the trajectories. DragDiffusion is inspired by the image editing method in (Shi et al., 2024), which utilizes feature maps extracted from the outputs of upsampling blocks to guide the generation process without feature alignment. MOFT instead utilizes feature maps derived from Content Correlation Removal proposed by Xiao et al. (2024a) known to encode motion information.
@@
Following prior works (Wu et al., 2024c; Zhou et al., 2024), we evaluate our method on the validation set of the VIPSeg dataset (Miao et al., 2022). We test on the same control regions and target trajectories as DragAnything, where the size of our bounding boxes is the same as the diameter of the circles in their work. For quantitative metrics, we report Frechet Inception Distance (FID) (Heusel et al., 2017) and Frechet Video Distance (FVD) (Unterthiner et al., 2018) to measure the visual quality, and ObjMC (Wu et al., 2024c) to measure the motion fidelity. ObjMC computes the average distance between generated and target trajectories, where Co-Tracker (Karaev et al., 2024) is used to estimate the trajectory of generated videos.